OpenStudy (anonymous):

Maria has three children. There is two years age difference between each child. The total ages of all three children is 36 years. Rosa is the oldest child. How old is Rosa? Let R = the age of Rosa. Which formula will calculate Rosa's age?

OpenStudy (anonymous):

\(R+(R-2)+((R-2)-2)=36\)

OpenStudy (apoorvk):

Let the age of Rosa be 'T' then age of next child = T-2 age of youngest child = 2 lesser than the prev. one = T-2 -2 = T-4 sum of ages = T + T -2 + T-4 = 3T-6 = 36 Solve and find 'T'.